OpenStudy (m):

Y = 3X - 2 Var(X) = .72 Var(Y) = 9Var(X) where did the 9 come from?

OpenStudy (anonymous):

Var(Y)=E((Y-E(Y))^2)=E((3X-2-(3E(X)-2))^2)=E((3X-3E(X))^2)= =E((3^2)(X-E(X))^2)=9E((X-E(X))^2)=9Var(X). Var(aX)=(a^2)Var(X)